OpenStudy (anonymous):

For which system of equations is (5, 3) the solution? A. 3x – 2y = 9 3x + 2y = 14 B. x – y = –2 4x – 3y = 11 C. –2x – y = –13 x + 2y = –11 D. 2x – y = 7 2x + 7y = 31

OpenStudy (uri):

@aye_itspaddie Solve each and see.

OpenStudy (uri):

Do you know how to solve?

OpenStudy (uri):

I'm a she. for the 10000000th time.

OpenStudy (uri):

@aye_itspaddie Solve each answer option to see if you get the same solution or not,If you don't know how to solve them just tell me.

OpenStudy (anonymous):

i dont know how to solve them i have the iQ of a grape.

OpenStudy (anonymous):

@uri

OpenStudy (uri):

Okay,I'll tell you.

OpenStudy (anonymous):

i just really neeed an answer

OpenStudy (anonymous):

i have 5 mins and 7 other questions

OpenStudy (uri):

3x – 2y = 9 3x + 2y = 14 ======= 6x=24 x=23/6 x=3.8 -2y+2y will be cancelled. No use to solve the whole question because x is 3.8,where as you want x to be 5.

OpenStudy (uri):

Well,I don't know the answer.....I just know how to solve.

OpenStudy (anonymous):

Put \(x=5\) and \(y=3\) in each equation and check which equation is satisfying these values.

OpenStudy (uri):

@waterineyes she wants only ansurs...

OpenStudy (anonymous):

Like I do for option D: \(2x - y = 7\) Put \(x=5\) and \(y=3\): \(2(5) - 3 = 7\) \(7=7\), so one equation is satisfied, check for others. :)

OpenStudy (anonymous):

Okay I got you Miss Uri. :)